Mercury Bag Company produces cases of grocery bags. The managers at Mercury are trying to develop budgets for the upcoming quarter. The following data have been gathered.

  

Projected sales in units

1,170

cases

Selling price per case

$

240

Inventory at the beginning of the quarter

150

cases

Target inventory at the end of the quarter

100

cases

Direct labor hours needed to produce one case

2

hours

Direct labor wages

$

10

per hour

Direct materials cost per case

$

8

Variable manufacturing overhead cost per case

$

6

Fixed overhead costs for the upcoming quarter

$

220,000

What are the budgeted sales in dollars?

What are the planned production amounts in cases?
Answer #1

Thanks for giving chance

For any query mention in comment section please

Answer:

Budgeted Sales in dollars $        280,800
Budgeted production of cases 1120 Cases
Budgeted Total manufacturing Cost $        258,080
Budgeted value of ending inventory $    23,042.86
Sales forecast
Budgeted sales (in cases) 1170
Selling price per case $                240
Budgeted sales (in dollars) = 1170 *240= $        280,800
Production schedule
Budgeted sales (cases) 1170
Target ending inventory 100
Cases budgeted to be available for sale 1270
Less: Beginning inventory 150
Planned production(cases) 1120
Calculation of Variable Manufacturing Cost per unit
Variable Manufacturing Cost
Direct Material $             8 Per Case
Direct Labor =2 hours per case x $10 per hour $          20 Per Case
Variable Overhead per unit $             6 Per Case
Total Variable Cost per case $          34 Per Case
Total Manufacturing cost budget
Per Case For 1120 Cases
Variable Manufacturing Cost
Direct Material 8 =1120*8 $               8,960
Direct Labor 20 =1120*20 $             22,400
Variable Overhead per unit 6 =1120*6 $               6,720
Total Variable Manufacturing Cost $             38,080
Fixed manufacturing overhead $           220,000
Total Manufacturing Cost of Finished Goods Manufactured $           258,080
Value of ending inventory = 258080 * 100/1120 $    23,042.86

Hit Thumbs up if satisfied

Have any query mention in comment section please

Thank you